OpenStudy (anonymous):

how to solve for x ? x-4/3=7/8

OpenStudy (lgbasallote):

im just writing the problem for the sake of \(\LaTeX\) \[\Large \frac{x-4}{3} = \frac 78\]

OpenStudy (anonymous):

mmmmmmm any idea how to solve for x?

OpenStudy (anonymous):

\[X-4/3=7/8\] cross multiply 8 to x-4, same with 7 and 3 and simplify the equation.

OpenStudy (lgbasallote):

yeah i have an idea but like i said i was just writing the problem hehe

OpenStudy (anonymous):

how do you multiply x-4 with 8? what would that become? thats what i dont get.

OpenStudy (krabsen):

8*(x-4)

OpenStudy (anonymous):

oh i see isee:)

OpenStudy (anonymous):

8x-32

OpenStudy (anonymous):

im guessing it wont end up going in evenly..

OpenStudy (anonymous):

8x-32=21 8x=53 x=53/8

OpenStudy (anonymous):

thankyou:)

OpenStudy (anonymous):

welcome